OpenStudy (anonymous):

PART:2 Consider the vectors u=(8,3) and v=(2,7), Determine proj v U.

OpenStudy (goformit100):

its a question of multiplication of scalar by vector

OpenStudy (anonymous):

Okay , can you help me with solving it?

OpenStudy (anonymous):

u.v/|v|=(8,3).(2,7)/sqrt18=8*2+3*7/sqrt18=16+21/sqrt18=37/sqrt18